# LatencyPrism: Online Non-intrusive Latency Sculpting for SLO-Guaranteed LLM Inference

### v1

@@ Abstract
We present LatencyPrism, the first zero-intrusion multi-platform latency sculpting system. It aims to break down the inference latency across pipeline, proactively alert on inference latency anomalies, and guarantee adherence to SLOs, all without requiring code modifications or service restarts. LatencyPrism has been deployed across thousands of XPUs for over six months. It enables low-overhead real-time monitoring at batch level with alerts triggered in milliseconds. This approach distinguishes between workload-driven latency variations and anomalies indicating underlying issues with an F1-score of 0.98. We also conduct extensive experiments and investigations into root cause analysis to demonstrate LatencyPrism’s capability.
@@ 1 Introduction
    Comprehensive Validation. We evaluate LatencyPrism against five categories of real-world scenarios, including computation contention, bandwidth saturation, and throttling. Experiments and deployment data show that LatencyPrism achieves an F1-score of 0.98 in anomaly detection. The high-fidelity context data effectively supports downstream root cause analysis and significantly accelerates manual troubleshooting.
@@ 4.2.2 Predictive Modeling and Feature Selection
4.2.3 Alert Triggering and Comprehensive Collection
@@ 5.1 Experimental Setup
We evaluated LatencyPrism on NVIDIA A100 GPUs using SGLang v0.5.4 and the Qwen3-32B model. The workload distribution spanned a wide range to mimic production diversity: batch sizes of 1–512, input lengths of 1–2048, and output lengths of 1–512.
5.2 Monitoring Subsystem Assessment
@@
6 Root Cause Localization
@@ 6.1 Metrics and Correlation
6.2 Suspicion Score Algorithm
@@ 7 Related Work
Reliability for Distributed AI Workloads. These systems target distributed training reliability. However, their localization precision is limited; for example, Mycroft [mycroft] cannot distinguish between unexpected performance issues and intentional overlap optimizations. Furthermore, Mycroft relies on instrumenting the NCCL library, imposing deployment and maintenance costs. Minder relies on metric thresholds, unable to distinguish root causes. Another example is XPUTimer [xputimer], which also requires code modification.
@@ 9 Conclusion
As Large Language Models transition from research artifacts to critical utility services, the operational focus must shift from training throughput to inference stability. We present LatencyPrism, the first zero-intrusion, full-stack performance tracking and anomaly detection system for production LLM inference. LatencyPrism solves the long-standing disconnection between upper-layer business logic and underlying physical resource states. Through large-scale deployment and controlled experiments, we validated its effectiveness: it maintains negligible overhead (<<0.5% CPU), achieves precise anomaly detection (F1=0.985) via dynamic baseline modeling, and provides rich, alignable trace data that accelerates root cause analysis. LatencyPrism demonstrates that building an intelligent, transparent, and efficient observability system is essential for guaranteeing service quality and unlocking computational potential in complex AI infrastructures.
@@ Acknowledgments
We provide the source code of core logic and the experimental results at https://github.com/kunluninsight/LatencyPrism.

### v2

@@ Abstract
We present LatencyPrism, the first zero-intrusion multi-platform latency sculpting system. It aims to break down the inference latency across pipeline, proactively alert on inference latency anomalies, and guarantee adherence to SLOs, all without requiring code modifications or service restarts. LatencyPrism has been deployed across thousands of XPUs for over six months. It enables low-overhead real-time monitoring at batch level with alerts triggered in milliseconds. This approach distinguishes between workload-driven latency variations and anomalies indicating underlying issues with an F1-score of 0.98. We also conduct extensive experiments and investigations into root cause analysis to demonstrate LatencyPrism’s capability. Furthermore, we introduce the first LLM anomaly simulation toolkit to facilitate future research in robust and predictable inference systems.
@@ 1 Introduction
    Comprehensive Validation. We evaluate LatencyPrism against several categories of real-world scenarios, including computation contention, bandwidth saturation, and throttling. To support this, we open-source the first comprehensive anomaly simulation suite specifically tailored for Large Model Distributed Inference. Experiments and deployment data show that LatencyPrism achieves an F1-score of 0.98 in anomaly detection. The high-fidelity context data effectively supports downstream root cause analysis and significantly accelerates manual troubleshooting.
@@ 4.2.2 Predictive Modeling and Feature Selection
4.2.3 Alert Trigger and Comprehensive Collection
@@
To evaluate LatencyPrism under diverse and realistic stress conditions, we developed an anomaly simulation suite. These simulations are organized according to the fundamental components of a modern, accelerator-equipped Von Neumann architecture, allowing us to systematically inject performance regressions by targeting specific hardware resources.
[…]
  • Processing Units: We simulate contention and performance degradation within the primary computational components of the system.
    • CPU (Central Processing Unit): To stress the CPU’s Arithmetic Logic Units (ALUs), we utilize AVX-512 saturation with dense floating-point loops. To simulate thermal throttling or power-saving modes, we enforce Dynamic Voltage and Frequency Scaling (DVFS) constraints using cpupower.
    • GPU (Graphics Processing Unit): We generate contention on the GPU’s Streaming Multiprocessors (SMs) by launching high-intensity General Matrix Multiply (GEMM) kernels. Additionally, we simulate power-limited or thermally constrained scenarios by locking the graphics and memory clocks to their minimum supported frequencies via nvidia-smi -lgc.
  • Memory System: We target the memory hierarchy to simulate Memory Thrashing and Fragmentation. We induce a state of high paging activity where the system excessively swaps data between RAM and disk, leading to severe performance degradation due to high I/O wait times. It can also break available memory into small, non-contiguous blocks after numerous allocation and deallocation cycles.
  • Interconnects and Data Bus: We create bottlenecks on the data pathways connecting the system’s components.
    • GPU-to-GPU Interconnect (NVLink): We saturate the high-speed NVLink bus through continuous peer-to-peer memory transfers, simulating heavy multi-GPU workloads.
    • CPU-to-GPU Interconnect (PCIe): We saturate the PCIe bus by programmatically disabling peer-to-peer (P2P) communication. This forces data transfers between GPUs to be relayed through host memory (Device $\rightarrow$ Host $\rightarrow$ Device), creating a significant bottleneck.
    • Bus Contention: We simulate contention from co-located devices, such as multiple GPUs and high-speed network interface cards (NICs) competing for limited PCIe lane bandwidth. This scenario mimics resource competition in densely populated server nodes, where one device’s high traffic can starve another.
[…]
@@ 5.1 Experimental Setup
Hardware and Models. Unless otherwise stated, we evaluated LatencyPrism on NVIDIA A100 GPUs using SGLang v0.5.4 and the Qwen3-32B model.
Data Scope and Granularity. Data was captured with all raw traces formatted for compatibility with standard visualization tools like Perfetto.
  • Sentinel Mode Evaluation: The workload distribution spanned a wide range to mimic production diversity: batch sizes of 1–512, input lengths of 1–2048, and output lengths of 1–512.
  • Deep Dive Mode Evaluation: We recorded high-density traces for each anomaly scenario. Each trace encompasses multiple concurrent threads and logs millions of fine-grained events, providing the statistical basis for our exploratory root cause localization.
5.2 Sentinel Mode Assessment
@@
5.3 Deep Dive Mode Assessment
@@ 5.3.1 Metrics and Correlation
5.3.2 Suspicion Score Algorithm
@@ 6 Related Work
Reliability for Distributed AI Workloads. These systems target distributed training reliability. However, their localization precision is limited; for example, Mycroft [mycroft] cannot distinguish between unexpected performance issues and intentional overlap optimizations. Furthermore, Mycroft relies on instrumenting the NCCL library, imposing deployment and maintenance costs. Minder [minder] relies on metric thresholds, unable to distinguish root causes. Another example is XPUTimer [xputimer], which also requires code modification.
@@ 8 Conclusion
As Large Language Models transition from research artifacts to critical utility services, the operational focus must shift from training throughput to inference stability. We present LatencyPrism, the first zero-intrusion, full-stack performance tracking and anomaly detection system for production LLM inference. LatencyPrism solves the long-standing disconnection between upper-layer business logic and underlying physical resource states. Through large-scale deployment and controlled experiments, we validated its effectiveness: it maintains negligible overhead (<<0.5% CPU), achieves precise anomaly detection (F1=0.985) via dynamic baseline modeling, and provides rich, alignable trace data that accelerates root cause analysis. LatencyPrism demonstrates that building an intelligent, transparent, and efficient observability system is essential for guaranteeing service quality and unlocking computational potential in complex AI infrastructures. Furthermore, we release the first open-source anomaly simulation suite, lowering the barrier for future research to investigate transient performance regressions in complex AI infrastructures.
@@ Acknowledgments
We provide the source code of core logic, anomaly injection scripts, and the experimental results at https://github.com/kunluninsight/LatencyPrism.